1 Evaluate ; dz,Cistheellipse + y2 = 1 c 2:+7:+3
Answer #1

Q. Answeron let fa)= 22 +72+3 & ginen inen C is an ellipse x² +42=1 now, Breaking fee), into partial fractione to (27+D (2+3)now o feidz = dz 22+72+3 С dz - (2Z FJ (213) =20 <dz -2 6 Idz c(27+1) 5, Z+ رف dz -6 I dz (A+12 Z+ Recall: - Cauchys theoremfa)=16 fel dz (7-00 21TU Ro Carne is an ellibre (. x² + y2 =) 4 Tunz how, I tlet I, - & 1 dz cl2th С Rez o Z=4 (20) He-va isClearly z =-3, Ų outside the Curve C. & hence is tic Cinside fonc) + Zt3 d auchys Whing Cauchys theorem Iz (2+3) how from